# Detecting and Quantifying Isolated Singularities over Discrete Valuation Rings

### v1

@@ 3 Power Series over a DVR
Since $\varphi$ is an isomorphism then $\mathfrak{m}_{V}=\langle\text{pr}\left(\varphi\left(x_{1}\right)\right),\dots,\text{pr}\left(\varphi\left(x_{n}\right)\right),\text{pr}\left(\varphi\left(y\right)\right)\rangle$. So we can set $H\left(z_{1},\dots,z_{n+1}\right)=\sum_{\alpha}u_{\alpha}z_{n+1}^{n_{\alpha}}\underline{z}^{\alpha}$ where $\underline{z}=\left(z_{1},\dots,z_{n}\right)$ together with $a_{i}=\text{pr}(\varphi(x_{i}))$ for $i=1,\dots,n$ and $a_{n+1}=\text{pr}(\varphi(y))$, and the result follows. ∎
The following proposition is a mixed characteristic version of Lemma 2.10 in [greuel2007introduction] and of Lemma 1.2.7 in [boubakri2009hypersurface].
@@
The following proposition shows that regular hypersurfaces are exactly of order $1$ and have a specific form up to equivalence (as a mixed characteristic version of Lemma 2.44 in [greuel2007introduction]).
@@
The following proposition is an analogue of Corollary 2.24 in [greuel2007introduction] and of Lemma 2.6 in [greuel1990simple].
@@
The following proposition is an analogue over $V[[\underline{x}]]$ of the splitting lemma over field (see, for example, Theorem 2.47 of [greuel2007introduction] and Lemma 3.9 of [greuel2016right]), and it is a generalized analogue of Lemma 3.1 in [carvajal2019covers]. For more information on the splitting lemma over fields, see [greuel2025splitting]. In addition, its proof is inspired by the proof of Theorem 3.7 in [svoray2025ade].
@@
Assume that $\kappa$ is quadratically closed and of characteristic different than $2$. Let $f\in V[[\underline{x}]]$ be of order $2$. Then there exists some $0\leq k\leq n$ and some $g\left(x_{k+1},\dots,x_{n}\right)\in\mathfrak{a}_{3}\subset V[[x_{k+1},\dots,x_{n}]]$ such that
@@
Now, by Definition 3.20 we can write $f_{1}=\vec{x}H\vec{x}^{\intercal}$ where $H=\operatorname{Hess}(f)$. Since $H$ is a symmetric matrix whose coefficients are either units or zero, then $f_{1}$ is a quadratic form, and so by Proposition 3.4 in Chapter I of [baeza2006quadratic] (noting that $2$ is invertible in $V$ since the characteristic of $\kappa$ is not $2$) there exists some invertible matrix $W$ and some diagonal matrix $D$ such that $H=WDW^{\intercal}$. Denote $\vec{w}=\vec{x}W=(w_{1},\dots,w_{n+1})$. Note that since $W$ is invertible then by Lemma 2.13 we can conclude that the $V$-homomorphism that sends $x_{i}\mapsto w_{i}$ for every $i\leq n$ and $y\mapsto w_{n+1}$ is a $V$-automorphism. In addition, up to rearranging the variables, we can denote the diagonal entries of $D$ by $u_{1},\dots,u_{k},u_{k+1},\dots,u_{n+1}$ such that $u_{1},\dots,u_{k}$ are units and $u_{k+1},\dots,u_{n+1}\in\mathfrak{n}_{V}$. Therefore we have that
@@
The following lemma tells us that the residual part of $f$ is well defined up to equivalence. It is a mixed characteristic version of part 2 of Lemma 9.2.10 in [de2013local] and of part 1 of Theorem 2.1 in [greuel2025splitting].
@@
Given $f,g\in\mathfrak{a}_{3}\subset V[[x_{1},\dots,x_{l},y]]$, we have that $f\left(x_{1},\dots,x_{l}\right)\overset{\mathcal{K}}{\sim}g\left(x_{1},\dots,x_{l}\right)$ if and only if $f\left(x_{1},\dots,x_{l}\right)+y^{2}+x_{l+1}^{2}+\dots+x_{n}^{2}\overset{\mathcal{K}}{\sim}g\left(x_{1},\dots,x_{l}\right)+y^{2}+x_{l+1}^{2}+\dots+x_{n}^{2}$ in $V[[\underline{x},y]]$.
@@
First, if $f\left(x_{1},\dots,x_{l}\right)\overset{\mathcal{K}}{\sim}g\left(x_{1},\dots,x_{l}\right)$ then there exists some unit $u\in V[[x_{1},\dots,x_{l},y]]$ and some automorphism $\varphi$ of $V[[x_{1},\dots,x_{l},y]]$ such that $f=u\varphi(g)$. So, by extending $\varphi$ to an automorphism $\psi$ of $V[[\underline{x},y]]$ by setting $\psi\left(x_{i}\right)=v^{-1}x_{i}$ for every $i>l$, where $v^{2}=u$, we can conclude that
@@
Now, write $\psi_{i}=u^{-0.5}\varphi_{i}=a_{i}+h_{i}$ where $a_{i}$ is a linear term over $\langle x_{1},\dots,x_{n},y\rangle$ and $h_{i}\in\mathfrak{n}_{V}^{2}$. Then we have that $\sum_{i=l+1}^{n+1}a_{i}^{2}-\sum_{i=l+1}^{n+1}x_{i}^{2}-y^{2}\in\mathfrak{n}_{V}^{2}$. In addition, we have that
$\sum_{i=l+1}^{n+1}\psi_{i}^{2}=\sum_{i=l+1}^{n+1}a_{i}^{2}+\sum_{i=l+1}^{n+1}h_{i}(2a_{i}+h_{i})$
Since $\operatorname{ord}_{\langle x_{1},\dots,x_{n},y\rangle}(a_{l+1})=\cdots=\operatorname{ord}_{\langle x_{1},\dots,x_{n},y\rangle}(a_{n+1})=1$, we have that $\sum_{i=1}^{n+1}a_{i}^{2}=\sum_{i=l+1}^{n}x_{i}^{2}+y^{2}$. Now, if we look at $F_{i}=2a_{i}+h_{i}$ (recalling that the characteristic of $\kappa$ is not $2$) then
@@
where $x_{n+1}=y$. Therefore, by the implicit function theorem and the inverse function theorem (Lemma 2.13), there exists some $\theta_{l+1},\dots,\theta_{n+1}\in V[[x_{1},\dots,x_{l}]]$ for which we have that $F_{i}\left(x_{1},\dots,x_{l},\theta_{l+1},\dots,\theta_{n+1}\right)=0$ for every $i$. For every $i\leq l$ define $\beta_{i}=\psi_{i}\left(x_{1},\dots,x_{l},\theta_{l+1},\dots,\theta_{n+1}\right)$. Then the map $\beta$ on $V[[x_{1},\dots,x_{l},y]]$ defined by $\beta\left(x_{i}\right)=\beta_{i}$ must be an automorphism. This gives us that the map defined by $\left(\psi_{1},\dots,\psi_{l},a_{l+1}+\frac{h_{l+1}}{2},\dots,a_{n+1}+\frac{h_{n+1}}{2}\right)$ is an automoprhism of $V[[\underline{x},y]]$. Thus, by the inverse function theorem (Lemma 2.13) we have that $\langle\psi_{1},\dots,\psi_{l},a_{l+1}+\frac{h_{l+1}}{2},\dots,a_{n+1}+\frac{h_{n+1}}{2}\rangle=\mathfrak{n}_{V}$. Therefore, for every $i\leq l$ there exists some $\alpha_{1,i},\dots,\alpha_{n+1,i}$ such that
@@
So $\langle\beta_{1},\dots,\beta_{l}\rangle=\langle x_{1},\dots,x_{l}\rangle$, and by the inverse function theorem (Lemma 2.13), we have that $\beta(f)=f\left(\beta_{1},\dots,\beta_{l}\right)=u\left(g\left(x_{1},\dots,x_{l}\right)\right)$ and the result follows. ∎
@@
and so after tensoring by $\frac{V[[\underline{x}]]}{\langle x_{l+1},\dots,x_{n}\rangle}$ we get an isomorphism of $V$-algebras $\frac{V[[x_{1},\dots,x_{l},y]]}{J(\tilde{f})}\cong\frac{V[[x_{1},\dots,x_{l},y]]}{J(\tilde{g})}$.
The following proposition gives us a classification of all $f\in V[[\underline{x}]]$ of full rank, which we can view as an analogue version of Morse’s lemma over $V[[\underline{x}]]$, as presented in Theorem 2.46 of Chapter I in [greuel2007introduction] (which in turn is an analytic version of Morse’s lemma, as presented in [morse1934calculus, varchenko1985singularities]). Note that it is also an analogue of Proposition 3.3. in [carvajal2019covers], which can be thought of as a mixed characteristic version of Morse’s lemma for surfaces, and its generalization in Proposition 3.14 of [svoray2025ade].
@@
    We can prove a variant of Proposition 3.27 based on Mather-Yau (Proposition 2.28). If we denote $f=\pi^{2}+x_{1}^{2}+\cdots+x_{n}^{2}$, then by Proposition 3.17 (or via similar computation to the one preformed in Example 2.2) we have that $f$ is 22-determined. Therefore, given some $g\in V[[\underline{x}]]$ with $J(\tilde{g})=\langle x_{1},\dots,x_{n},y\rangle$, we have that $\frac{V[[\underline{x},y]]}{J(\tilde{f})}\cong\frac{V[[\underline{x},y]]}{J(\tilde{g})}$ and so $f\sim g$.
@@
    Inspired by Proposition 3.27 one might want to preform an ADEADE-like classification up to $\sim$ over $V[[\underline{x}]]$ based upon the the rank of order $2$ elements (as in Section 2.4 of chapter I in [greuel2007introduction] over $\mathbb{C}$, as in [greuel1990simple, Nguyen2013classification] over algebraically closed field of positive characteristic, or as in [carvajal2019covers, svoray2025ade] in the mixed characteristic case). Yet, in Proposition 3.22 we get that the residual part is a member of $\mathfrak{a}_{3}$, and therefore it is unclear if it would be $\overset{\mathcal{K}}{\sim}$ to some $\tilde{h}\in V[[\underline{x}]]$, which would give us an equivalence $\sim$ over $V[[\underline{x}]]$.
@@ 4 Isolated Singularities over a DVR
In this section we show how the work of Hochster and Jefferies in [hochster2021jacobian], Saito in [saito2022frobenius], and KC in [kc2024singular] can inspire a similar construction to our Tjurina number that detects singularities over unramified and ramified DVRs, as opposed to Proposition 3.6.
We start with the unramified case. Specifically, we assume that $(V,\mathfrak{m}_{V},\kappa)$ is a complete DVR of mixed characteristic $(0,p)$ whose uniformizer is $p$ (i.e. $\mathfrak{m}_{V}=\langle p\rangle$) and $\kappa$ is algebraically closed. In this case we want to quantify when $f\in V[[\underline{x}]]$ defines an isolated singularity. We first review a few facts about $p$-derivation, as described in [hochster2021jacobian].
@@
By the Cohen structure theorem (see [cohen1946structure]), every complete DVR with uniformizer $p$ and an algebraically closed (therefore perfect) quotient field $\kappa$ must be isomorphic to the ring of Witt vectors over $\kappa$. For more about this connection between Witt vectors and Cohen rings, see [anscombe2022model] or Section 6 of [hazewinkel2009witt], and for a deeper review of Witt vectors, see [rabinoff2014theory] or Chapter II, Section 6 of [serre1979local].
@@
Note that $C_{p}\left(X_{1},X_{2}\right)\in\langle X_{1}X_{2}\rangle$ and in general we have that $C_{p}\left(X_{1},\dots,X_{n}\right)\in\langle X_{1},\dots,X_{n}\rangle^{2}$.
The following definition is a combination of Definition 2.2 and Definition 2.5 in [hochster2021jacobian].
@@
The notion of $p$-derivations goes back to the works of Buium in [buium1995differential] and Jóyal in [joyal1985delta]. In addition, Jóyal proved that the forgetful functor from the category of rings with $p$-derivations to the category of rings is the left adjoint of the Witt vector functor (with the natural $p$-derivation). One can view $p$-derivations as "derivations with respect to pp", and play an analogous role to derivations, as we see with the mixed characteristic Jacobian criterion in Theorem 4.8 or with the mixed characteristic Zariski-Nagata Theorem in [de2020zariski, de2025differential].
The following lemma summarizes a few results about $p$-derivations that are presented in Section 2 of [hochster2021jacobian].
@@
    For more on the construction of $\operatorname{Frob}$ as the lift of the Frobenius map and for its properties, see [davis2014witt].
@@
Note that by item 2 of Lemma 4.4 it is enough to prove the result for $f\in V[[\underline{x}]]$ that are polynomial, as we can approximate $\delta_{\underline{a}}(f)$ by looking at the $\delta_{\underline{a}}$ of the finite Taylor expansions of $f$. In addition, since sum and product are continuous with respect to the $\mathfrak{m}_{V}$-adic topology, it is enough to prove this result for monomials in $p,x_{1},\dots,x_{n}$, which we prove via induction. The map $\delta\mapsto\delta(x_{i})$ is continuous since it is simply the projection from the product topology. In addition, note that by Lemma 4.4 we can conclude that $\delta(p^{j})=\frac{p^{j}-p^{jp}}{p}$. Now, if $p^{j}\underline{x}^{\alpha}$ is a monomial (with $\alpha\neq\underline{0}$), we can write it as $p^{j}\underline{x}^{\alpha}=x_{l}\underline{x}^{\beta}$, and so $\delta(p^{j}\underline{x}^{\alpha})=\delta(x_{l}p^{j}\underline{x}^{\beta})=x_{l}^{p}\delta(p^{j}\underline{x}^{\beta})+p^{jp}\underline{x}^{p\beta}\delta(x_{l})+p\delta(x_{l})\delta(p^{j}\underline{x}^{\beta})$, which is continuous as a combination of continuous functions. ∎
@@
We now recall the main result of [hochster2021jacobian] (in our case and using our notational convention) which we use in order to understanding and quantify isolated singularities using $p$-derivations.
Theorem 4.8 (Theorem 4.9. in [hochster2021jacobian]).
@@
The second part Theorem 4.8 was first presented in [nagata1958general], and is in fact true for every mixed-characteristic complete DVR.
@@
Given some $f\in\mathfrak{m}_{V}$, we have that $\frac{V[[\underline{x}]]}{\langle f\rangle}$ is regular if and only if the length of $\frac{V[[\underline{x}]]}{\langle f,\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle}$ over $V[[\underline{x}]]$ is zero, for any $p$-derivation $\delta$.
@@
Since $V[[\underline{x}]]$ is local then we have that $\frac{V[[\underline{x}]]}{\langle f\rangle}=\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{m}_{V}}$. Yet, since $f,p\in\mathfrak{m}_{V}$ then by Theorem 4.8, $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{m}_{V}}$ is regular if and only if for every $\delta$ we have that $\langle\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle\not\subset\mathfrak{m}_{V}$. But, $\langle\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle\not\subset\mathfrak{m}_{V}$ if and only if $\langle\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle=V[[\underline{x}]]$, and since 0 is the only module that has length zero, the result follows. ∎
@@
If $f$ defines an isolated singularity, by Theorem 4.8 for every $\delta$ and for every $\mathfrak{q}\neq\mathfrak{m}_{V}$ we have that $J_{\delta}\left(f\right)+\langle f\rangle\not\subset\mathfrak{q}$ if $p\in\mathfrak{q}$ and $J\left(f\right)\not\subset\mathfrak{q}$ if $p\notin\mathfrak{q}$, and so $\tau(f,\delta)<\infty$. In addition, $p\in\sqrt{J\left(f\right)}$ since if $f$ defines an isolated singularity, then in particular, $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)[\frac{1}{p}]$ is regular, and so as its completion with respect to $x_{1},\dots,x_{n}$, which is isomorphic to $\frac{L[[\underline{x}]]}{\langle f\rangle}$, where $L$ is the fraction field of $V$. Since $L$ is a field of characteristic zero, then by Theorem 5.1.7. of [boubakri2009hypersurface] together with Lemma 2.44 in [greuel2007introduction] we must have that $1\in J\left(f\right)\otimes_{V}L[[\underline{x}]]$, and so $1\in J\left(f\right)[\frac{1}{p}]$.

### v2

@@ 3 Power Series over a DVR
Since $\varphi$ is an isomorphism then we must have that that the ideal defined by $\langle\text{pr}\left(\varphi\left(x_{1}\right)\right),\dots,\text{pr}\left(\varphi\left(x_{n}\right)\right),\text{pr}\left(\varphi\left(y\right)\right)\rangle$ is exactly $\mathfrak{m}_{V}$. Therefore we can set $H\left(z_{1},\dots,z_{n+1}\right)=\sum_{\alpha}u_{\alpha}z_{n+1}^{n_{\alpha}}\underline{z}^{\alpha}$ where $\underline{z}=\left(z_{1},\dots,z_{n}\right)$ together with $a_{i}=\text{pr}(\varphi(x_{i}))$ for $i=1,\dots,n$ and $a_{n+1}=\text{pr}(\varphi(y))$, and the result follows. ∎
The following proposition is a mixed characteristic version of Lemma 2.10 in [GLS07] and of Lemma 1.2.7 in [BOU09].
@@
The following proposition shows that regular hypersurfaces are exactly of order $1$ and have a specific form up to equivalence (as a mixed characteristic version of Lemma 2.44 in [GLS07]).
@@
The following proposition is an analogue of Corollary 2.24 in [GLS07] and of Lemma 2.6 in [GK90].
@@
The following proposition is an analogue over $V[[\underline{x}]]$ of the splitting lemma over field (see, for example, Theorem 2.47 of [GLS07] and Lemma 3.9 of [GN16]), and it is a generalized analogue of Lemma 3.1 in [CMP+21]. For more information on the splitting lemma over fields, see [GP25]. In addition, its proof is inspired by the proof of Theorem 3.7 in [SVO25].
@@
Assume that the field $\kappa$ is quadratically closed and of characteristic different than $2$. Let $f\in V[[\underline{x}]]$ be of order $2$. Then there exists some $0\leq k\leq n$ and some $g\left(x_{k+1},\dots,x_{n}\right)\in\mathfrak{a}_{3}\subset V[[x_{k+1},\dots,x_{n}]]$ such that
@@
Now, by Definition 3.20 we can write $f_{1}=\vec{x}H\vec{x}^{\intercal}$ where $H=\operatorname{Hess}(f)$. Since $H$ is a symmetric matrix whose coefficients are either units or zero, then $f_{1}$ is a quadratic form, and so by Proposition 3.4 in Chapter I of [BAE06] (noting that $2$ is invertible in $V$ since the characteristic of $\kappa$ is not $2$) there exists some invertible matrix $W$ and some diagonal matrix $D$ such that $H=WDW^{\intercal}$. Denote $\vec{w}=\vec{x}W=(w_{1},\dots,w_{n+1})$. Note that since $W$ is invertible then by Lemma 2.13 we can conclude that the $V$-homomorphism that sends $x_{i}\mapsto w_{i}$ for every $i\leq n$ and $y\mapsto w_{n+1}$ is a $V$-automorphism. In addition, up to rearranging the variables, we can denote the diagonal entries of $D$ by $u_{1},\dots,u_{k},u_{k+1},\dots,u_{n+1}$ such that $u_{1},\dots,u_{k}$ are units and $u_{k+1},\dots,u_{n+1}\in\mathfrak{n}_{V}$. Therefore we have that
@@
The following lemma tells us that the residual part of $f$ is well defined up to equivalence. It is a mixed characteristic version of part 2 of Lemma 9.2.10 in [DP13] and of part 1 of Theorem 2.1 in [GP25].
@@
Let $f,g\in\mathfrak{a}_{3}\subset V[[x_{1},\dots,x_{l},y]]$. Then we have $f\left(x_{1},\dots,x_{l}\right)\overset{\mathcal{K}}{\sim}g\left(x_{1},\dots,x_{l}\right)$ in $V[[x_{1},\dots,x_{l}]]$ if and only if in $V[[\underline{x},y]]$ we have the equivalence
$f\left(x_{1},\dots,x_{l}\right)+y^{2}+x_{l+1}^{2}+\dots+x_{n}^{2}\overset{\mathcal{K}}{\sim}g\left(x_{1},\dots,x_{l}\right)+y^{2}+x_{l+1}^{2}+\dots+x_{n}^{2}.$
@@
First, if $f\left(x_{1},\dots,x_{l}\right)\overset{\mathcal{K}}{\sim}g\left(x_{1},\dots,x_{l}\right)$ then we can find some automorphism $\varphi$ of $V[[x_{1},\dots,x_{l},y]]$ and some unit $u\in V[[x_{1},\dots,x_{l},y]]$ such that $f=u\varphi(g)$. So, by extending $\varphi$ to an automorphism $\psi$ of $V[[\underline{x},y]]$ by setting $\psi\left(x_{i}\right)=v^{-1}x_{i}$ for every $i>l$, where $v^{2}=u$, we can conclude that
@@
Now, write $\psi_{i}=u^{-0.5}\varphi_{i}=a_{i}+h_{i}$ where $a_{i}$ is a linear term over $\langle x_{1},\dots,x_{n},y\rangle$ and $h_{i}\in\mathfrak{n}_{V}^{2}$. Then we have that
$\sum_{i=l+1}^{n+1}a_{i}^{2}-\sum_{i=l+1}^{n+1}x_{i}^{2}-y^{2}\in\mathfrak{n}_{V}^{2}.$
In addition, we have that
$\sum_{i=l+1}^{n+1}\psi_{i}^{2}=\sum_{i=l+1}^{n+1}a_{i}^{2}+\sum_{i=l+1}^{n+1}h_{i}(2a_{i}+h_{i}),$
and since $\operatorname{ord}_{\langle x_{1},\dots,x_{n},y\rangle}(a_{l+1})=\cdots=\operatorname{ord}_{\langle x_{1},\dots,x_{n},y\rangle}(a_{n+1})=1$, we can conclude that $\sum_{i=1}^{n+1}a_{i}^{2}=\sum_{i=l+1}^{n}x_{i}^{2}+y^{2}$. Now, if we look at $F_{i}=2a_{i}+h_{i}$ (recalling that the characteristic of $\kappa$ is not $2$) then
@@
where $x_{n+1}=y$. Therefore, by the implicit function theorem and the inverse function theorem (Lemma 2.13), there exists some $\theta_{l+1},\dots,\theta_{n+1}\in V[[x_{1},\dots,x_{l}]]$ for which we have that $F_{i}\left(x_{1},\dots,x_{l},\theta_{l+1},\dots,\theta_{n+1}\right)=0$ for every $i$. For every $i\leq l$ define $\beta_{i}=\psi_{i}\left(x_{1},\dots,x_{l},\theta_{l+1},\dots,\theta_{n+1}\right)$. Then the map $\beta$ on $V[[x_{1},\dots,x_{l},y]]$ defined by $\beta\left(x_{i}\right)=\beta_{i}$ must be an automorphism. This tells us that $\left(\psi_{1},\dots,\psi_{l},a_{l+1}+\frac{h_{l+1}}{2},\dots,a_{n+1}+\frac{h_{n+1}}{2}\right)$ defines a $V$-automoprhism of $V[[\underline{x},y]]$. Thus, by the inverse function theorem (Lemma 2.13) we have that $\langle\psi_{1},\dots,\psi_{l},a_{l+1}+\frac{h_{l+1}}{2},\dots,a_{n+1}+\frac{h_{n+1}}{2}\rangle=\mathfrak{n}_{V}$. Therefore, for every $i\leq l$ there exists some $\alpha_{1,i},\dots,\alpha_{n+1,i}$ such that
@@
Therefore we have that $\langle\beta_{1},\dots,\beta_{l}\rangle=\langle x_{1},\dots,x_{l}\rangle$, and by the inverse function theorem (Lemma 2.13), we can conclude that $\beta(f)=f\left(\beta_{1},\dots,\beta_{l}\right)=u\left(g\left(x_{1},\dots,x_{l}\right)\right)$ and the result follows. ∎
@@
and so after tensoring by $\frac{V[[\underline{x}]]}{\langle x_{l+1},\dots,x_{n}\rangle}$ we can conclude that we get an isomorphism of $V$-algebras $\frac{V[[x_{1},\dots,x_{l},y]]}{J(\tilde{f})}\cong\frac{V[[x_{1},\dots,x_{l},y]]}{J(\tilde{g})}$.
The following proposition gives us a classification of all $f\in V[[\underline{x}]]$ of full rank, which we can view as an analogue version of Morse’s lemma over $V[[\underline{x}]]$, as presented in Theorem 2.46 of Chapter I in [GLS07] (which in turn is an analytic version of Morse’s lemma, as presented in [MOR34, VG85]). Note that it is also an analogue of Proposition 3.3. in [CMP+21], which can be thought of as a mixed characteristic version of Morse’s lemma for surfaces, and its generalization in Proposition 3.14 of [SVO25].
@@
    We can prove a variant of Proposition 3.27 based on Mather-Yau (as presented in Proposition 2.33). If we denote $f=\pi^{2}+x_{1}^{2}+\cdots+x_{n}^{2}$, then by Proposition 3.17 (or via similar computation to the one preformed in Example 2.2) we have that $f$ is 22-determined. Therefore, given some $g\in V[[\underline{x}]]$ with $J(\tilde{g})=\langle x_{1},\dots,x_{n},y\rangle$, we have that $\frac{V[[\underline{x},y]]}{J(\tilde{f})}\cong\frac{V[[\underline{x},y]]}{J(\tilde{g})}$ and so $f\sim g$.
@@
    Inspired by Proposition 3.27 one might want to preform an ADEADE-like classification up to $\sim$ over $V[[\underline{x}]]$ based upon the the rank of order $2$ elements (as in Section 2.4 of chapter I in [GLS07] over $\mathbb{C}$, as in [GK90, NGU13] over algebraically closed field of positive characteristic, or as in [CMP+21, SVO25] in the mixed characteristic case). Yet, in Proposition 3.22 we get that the residual part is a member of $\mathfrak{a}_{3}$, and therefore it is unclear if it would be $\overset{\mathcal{K}}{\sim}$ to some $\tilde{h}\in V[[\underline{x}]]$, which would give us an equivalence $\sim$ over $V[[\underline{x}]]$.
@@ 4 Isolated Singularities over a DVR
In this section we show how the work of Hochster and Jefferies in [HJ21], Saito in [SAI22], and KC in [KC24] can inspire a similar construction to our Tjurina number that detects singularities over unramified and ramified DVRs, as opposed to Proposition 3.6.
We start with the unramified case. Specifically, we assume that $(V,\mathfrak{m}_{V},\kappa)$ is a complete DVR of mixed characteristic $(0,p)$ whose uniformizer is $p$ (i.e. $\mathfrak{m}_{V}=\langle p\rangle$) and $\kappa$ is algebraically closed. In this case we want to quantify when $f\in V[[\underline{x}]]$ defines an isolated singularity. We first review a few facts about $p$-derivation, as described in [HJ21].
@@
By the Cohen structure theorem (see [COH46]), every complete DVR with uniformizer $p$ and an algebraically closed (therefore perfect) quotient field $\kappa$ must be isomorphic to the ring of Witt vectors over $\kappa$. For more about this connection between Witt vectors and Cohen rings, see [AJ22] or Section 6 of [HAZ09], and for a deeper review of Witt vectors, see [RAB14] or Chapter II, Section 6 of [SER79].
@@
Note that $C_{p}\left(X_{1},X_{2}\right)\in\langle X_{1}X_{2}\rangle$ and, in general, for every integer $n$ we have that $C_{p}\left(X_{1},\dots,X_{n}\right)\in\langle X_{1},\dots,X_{n}\rangle^{2}$.
The following definition is a combination of Definition 2.2 and Definition 2.5 in [HJ21].
@@
The notion of $p$-derivations goes back to the works of Buium in [BUI95] and Jóyal in [JOY85]. In addition, Jóyal proved that the forgetful functor from the category of rings with $p$-derivations to the category of rings is the left adjoint of the Witt vector functor (with the natural $p$-derivation). One can view $p$-derivations as "derivations with respect to pp", and play an analogous role to derivations, as we see with the mixed characteristic Jacobian criterion in Theorem 4.8 or with the mixed characteristic Zariski-Nagata Theorem in [DGJ20, DGJ25].
The following lemma summarizes a few results about $p$-derivations that are presented in Section 2 of [HJ21].
@@
    For more on the construction of $\operatorname{Frob}$ as the lift of the Frobenius map and for its properties, see [DK14].
@@
Note that by item 2 of Lemma 4.4 it is enough to prove the result for $f\in V[[\underline{x}]]$ that are polynomial, as we can approximate $\delta_{\underline{a}}(f)$ by looking at the $\delta_{\underline{a}}$ of the finite Taylor expansions of $f$. In addition, since sum and product are continuous with respect to the $\mathfrak{m}_{V}$-adic topology, it is enough to prove this result for monomials in $p,x_{1},\dots,x_{n}$, which we prove via induction. The map $\delta\mapsto\delta(x_{i})$ is continuous since it is simply the projection from the product topology. In addition, note that by Lemma 4.4 we can conclude that $\delta(p^{j})=\frac{p^{j}-p^{jp}}{p}$. Now, if $p^{j}\underline{x}^{\alpha}$ is a monomial (with $\alpha\neq\underline{0}$), we can write it as $p^{j}\underline{x}^{\alpha}=x_{l}\underline{x}^{\beta}$, and so
$\delta(p^{j}\underline{x}^{\alpha})=\delta(x_{l}p^{j}\underline{x}^{\beta})=x_{l}^{p}\delta(p^{j}\underline{x}^{\beta})+p^{jp}\underline{x}^{p\beta}\delta(x_{l})+p\delta(x_{l})\delta(p^{j}\underline{x}^{\beta}),$
which is continuous as a combination of continuous functions. ∎
@@
We now recall the main result of [HJ21] (in our case and using our notational convention) which we use in order to understanding and quantify isolated singularities using $p$-derivations.
Theorem 4.8 (Theorem 4.9. in [HJ21]).
@@
The second part Theorem 4.8 was first presented in [NAG58], and is in fact true for every mixed-characteristic complete DVR.
@@
Given some $f\in\mathfrak{m}_{V}$, we have that $\frac{V[[\underline{x}]]}{\langle f\rangle}$ is regular if and only if for any $p$-derivation $\delta$, the length of $\frac{V[[\underline{x}]]}{\langle f,\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle}$ over $V[[\underline{x}]]$ is zero.
@@
Since $V[[\underline{x}]]$ is local then we have that $\frac{V[[\underline{x}]]}{\langle f\rangle}=\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{m}_{V}}$. Yet, since $f,p\in\mathfrak{m}_{V}$ then by Theorem 4.8, $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)_{\mathfrak{m}_{V}}$ is regular if and only if for every $\delta$ we have that $\langle\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle\not\subset\mathfrak{m}_{V}$. But we can conclude that $\langle\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle\not\subset\mathfrak{m}_{V}$ if and only if $\langle\left(\partial_{1}f\right)^{p},\dots,\left(\partial_{n}f\right)^{p},\delta\left(f\right)\rangle=V[[\underline{x}]]$, and since 0 is the only module that has length zero, the result follows. ∎
@@
If $f$ defines an isolated singularity, by Theorem 4.8 for every $\delta$ and for every $\mathfrak{q}\neq\mathfrak{m}_{V}$ we have that $J_{\delta}\left(f\right)+\langle f\rangle\not\subset\mathfrak{q}$ if $p\in\mathfrak{q}$ and $J\left(f\right)\not\subset\mathfrak{q}$ if $p\notin\mathfrak{q}$, and so $\tau(f,\delta)<\infty$. In addition, $p\in\sqrt{J\left(f\right)}$ since if $f$ defines an isolated singularity, then in particular, $\left(\frac{V[[\underline{x}]]}{\langle f\rangle}\right)[\frac{1}{p}]$ is regular, and so as its completion with respect to $x_{1},\dots,x_{n}$, which is isomorphic to $\frac{L[[\underline{x}]]}{\langle f\rangle}$, where $L$ is the fraction field of $V$. Since $L$ is a field of characteristic zero, then by Theorem 5.1.7. of [BOU09] together with Lemma 2.44 in [GLS07] we must have that $1\in J\left(f\right)\otimes_{V}L[[\underline{x}]]$, and so $1\in J\left(f\right)[\frac{1}{p}]$.